# ERATTA: Extreme RAG for enterprise-Table To Answers with Large Language Models

### v3

@@ Abstract
Large language models (LLMs) with retrieval augmented-generation (RAG) have been the optimal choice for scalable generative AI solutions in the recent past. Although RAG implemented with AI agents (agentic-RAG) has been recently popularized, its suffers from unstable cost and unreliable performances for Enterprise-level data-practices. Most existing use-cases that incorporate RAG with LLMs have been either generic or extremely domain specific, thereby questioning the scalability and generalizability of RAG-LLM approaches. In this work, we propose a unique LLM-based system where multiple LLMs can be invoked to enable data authentication, user-query routing, data-retrieval and custom prompting for question-answering capabilities from Enterprise-data tables. The source tables here are highly fluctuating and large in size and the proposed framework enables structured responses in under 10 seconds per query. Additionally, we propose a five metric scoring module that detects and reports hallucinations in the LLM responses. Our proposed system and scoring metrics achieve $>90\%$ confidence scores across hundreds of user queries in the sustainability, financial health and social media domains. Extensions to the proposed extreme RAG architectures can enable heterogeneous source querying using LLMs.
@@ I Introduction
Generative AI solutions using Large Language Models (LLMs) have been extremely popular since the launch of Open AI’s chatGPT-3 in late 2022 [1]. Most LLM solutions today can typically be categorized as Gen-AI applications, such as a chatbot-like experience with a virtual assistant for code, language and content generation [2]; or Gen-AI agents that are capable of automated orchestration to perform specific tasks such as booking tickets, writing blogs, articles and generating software automatically [3]. Most such solutions implement the retrieval augmented-generation (RAG) approach, wherein the most relevant data sources are first identified followed by the knowledge-graph (KG) method to isolate only relevant data entities that are collated with natural language instructions and answering guidelines and sent to the LLM to receive the required responses [4]. Although agentic-RAG approaches have been found to enhance the “knowledge-quality” of retrieved content instead of relying on LLM memory, these approaches can be costly and time-intensive owing to the web-search and KG method. Thus, for Enterprise-level data products involving large numbers of tabular rows and columns, agentic-RAG can prove impractical for question-answering for a large number of users/groups. In this paper, we present a novel system architecture that maximizes the RAG impact through multiple LLM calls for tasks such as user authentication, query routing and natural language question-answering while ensuring security/access requirements and accuracy standards by flagging hallucinations.
@@
Figure 1: The proposed extreme-RAG system architecture that invokes 3-4 LLM prompts per user query to authenticate, route, extract smaller tabular data and fetch the required response from fast-varying and large tabular data sources.
@@ V Conclusion
In this work, we propose a novel multi-LLM system that enables scalable tables-to-answers for Enterprise data tables with an averaged accuracy of $>90\%$ in under 10 seconds per user-query. The proposed process of dividing a user-query for query authentication, query understanding and routing, sub-tablular retrieval followed by answer generation is more stable and faster when compared to agentic-RAG approaches, which makes it a better fit for numeric-intensive use cases such as Sales, Financial and Sustainability domains. Future works can be directed towards extending RAG approaches to heterogeneous data sources for automation and insight generation tasks.

### v4

@@ Abstract
Large language models (LLMs) with retrieval augmented-generation (RAG) have been the optimal choice for scalable generative AI solutions in the recent past. Although RAG implemented with AI agents (agentic-RAG) has been recently popularized, its suffers from unstable cost and unreliable performances for Enterprise-level data-practices. Most existing use-cases that incorporate RAG with LLMs have been either generic or extremely domain specific, thereby questioning the scalability and generalizability of RAG-LLM approaches. In this work, we propose a unique LLM-based system where multiple LLMs can be invoked to enable data authentication, user-query routing, data-retrieval and custom prompting for question-answering capabilities from Enterprise-level data tables on sustainability. The source tables here are highly fluctuating and large in size storing carbon footprint, energy and water usage at buildings in regional levels globally and the proposed framework enables structured responses in under 10 seconds per query. Additionally, we propose a five metric scoring module that detects and reports hallucinations in the LLM responses. Our proposed system and scoring metrics achieve $>90\%$ confidence scores across hundreds of user queries in the sustainability, financial health and social media domains. Extensions to the proposed extreme RAG architectures can enable heterogeneous source querying using LLMs.
@@ I Introduction
Generative AI solutions using Large Language Models (LLMs) have been extremely popular since the launch of Open AI’s chatGPT-3 in late 2022 [1]. Most LLM solutions today can typically be categorized as Gen-AI applications, such as a chatbot-like experience with a virtual assistant for code, language and content generation [2]; or Gen-AI agents that are capable of automated orchestration to perform specific tasks such as booking tickets, writing blogs, articles and generating software automatically [3]. Most such solutions implement the retrieval augmented-generation (RAG) approach, wherein the most relevant data sources are first identified followed by the knowledge-graph (KG) method to isolate only relevant data entities that are collated with natural language instructions and answering guidelines and sent to the LLM to receive the required responses [4]. Although agentic-RAG approaches have been found to enhance the “knowledge-quality” of retrieved content instead of relying on LLM memory, these approaches can be costly and time-intensive owing to the web-search and KG method. Thus, for Enterprise-level data products involving large numbers of tabular rows and columns, agentic-RAG can prove impractical for question-answering for a large number of users/groups. In this paper, we present a novel system architecture that maximizes the RAG impact through multiple LLM calls for tasks such as user authentication, query routing and natural language question-answering while ensuring security/access requirements and accuracy standards by flagging hallucinations for critical domains such as sustainability.
@@
Figure 1: The proposed extreme-RAG system architecture that invokes 3-4 LLM prompts per user query to authenticate, route, extract smaller tabular data and fetch the required response from fast-varying and large tabular sustainability data sources.
@@ V Conclusion
In this work, we propose a novel multi-LLM system that enables scalable tables-to-answers for Enterprise-level sustainability data tables with an averaged accuracy of $>90\%$ in under 10 seconds per user-query. The proposed process of dividing a user-query for query authentication, query understanding and routing, sub-tablular retrieval followed by answer generation is more stable and faster when compared to agentic-RAG approaches, which makes it a better fit for numeric-intensive use cases such as Sales, Financial and Sustainability domains. Future works can be directed towards extending RAG approaches to heterogeneous data sources for automation and insight generation tasks.